2. A car travels form rest and has a constant acceleration of 3 ft/s^2. Find the average velocity during the first 10s of motion.

1
Expert's answer
2022-06-13T15:34:13-0400

Given:

"a=3\:\rm ft/s^2"

"v_0=0\:\rm ft/s"

"t=10\:\rm s"


The distance traveled by a car during the first 10s of motion

"d=v_0t+at^2/2=3*10^2/2=150\;\rm ft"

The average velocity during this period of motion

"v_{ave}=\frac{d}{t}=\frac{150}{10}=15\:\rm ft/s"